# Mean-Field Microcanonical Gradient Descent

### v1

@@ Abstract
Generative Models, Energy-Based Models, Microcanonical Models, Regularization, Entropy, Time Series
@@ 1 Introduction
The defining characteristic of a well-behaved generative model is the balance between its ability to, on the one hand produce samples that are typical of the training data, while on the other hand having a significant amount of diversity within its samples. For example, a generative adversarial network (GAN) which has suffered mode collapse could produce great samples within one mode but would then only produce samples within that one mode. On one extreme, we therefore have a model which exactly reproduces the empirical distribution of the training set, while on the other extreme, we have a model which produces samples of Gaussian white noise regardless of the training data. Formally, we can view this in terms of the reverse Kullback–Leibler (KL) divergence (Papamakarios et al., 2021) of the generative model $q$ with respect to the true distribution $p$:
$\mathcal{D}_{\text{KL}}(q\parallel p)=-H(q)-\mathbb{E}_{q}[\log p(X)],$
@@
One popular family of generative models is that of the energy-based model (Geman & Geman, 1984), also known as a canonical or macrocanonical model (Jaynes, 1957), which takes an energy vector $\Phi:\mathcal{X}\to\mathbb{R}^{K}$ defined on the sample space $\mathcal{X}$ and produces samples from the distribution that maximizes the entropy given the constraint $\mathbb{E}_{q}[\Phi(X)]=\alpha$, fixing the expected energy vector to some point $\alpha\in\mathbb{R}^{K}$. A macrocanonical model can in many cases be approximated by a microcanonical model, which similarly maximizes entropy, but over the set
$\Omega_{\varepsilon}=\{x\in\mathcal{X}:\|\Phi(x)-\alpha\|\leq\varepsilon\},$ (1)
for some $\varepsilon>0$ (Lanford, 1975; Ellis et al., 2000; Touchette, 2015). This $\Omega_{\varepsilon}$ is known as the microcanonical set of width $\varepsilon$. Without loss of generality, we shall assume that $\alpha=0$ (otherwise, we simply redefine $\Phi(x)$ to be $\Phi(x)-\alpha$). This microcanonical approximation is often useful when it is reasonable to assume that the energy vector concentrates in a region of high probability. For example, in the context of stationary time series of sufficiently long duration $d$, this amounts to working with energy vectors defined as the time average of time-shift equivariant potentials.
@@
While conceptually simple and theoretically useful, both macrocanonical and microcanonical models suffer from difficulties in sampling for high-dimensional sample spaces. To mitigate this, the microcanonical gradient descent model (MGDM) was introduced by Bruna & Mallat (2019) as an approximation of the microcanonical model which is easier to sample from. MGDM is defined as the pushforward of Gaussian white noise by way of a sequence of gradient descent steps that seek to minimize the objective
$L(x)=\frac{1}{2}\|\Phi(x)\|^{2}.$ (2)
@@
Thus samples from the MGDM are generated by sampling $x_{0}$ from $\mathcal{N}(0,\sigma^{2}\text{I}_{d})$ for some initial variance $\sigma^{2}$ and updating the sample using
$x_{t+1}=g(x_{t})\coloneqq x_{t}-\gamma J_{\Phi}^{\top}(x_{t})\Phi(x_{t}),$ (3)
@@
where $\gamma$ is the step size and $J_{\Phi}$ is the Jacobian of $\Phi$. This is typically iterated for a fixed number of steps $T$ or until $x_{t}$ reaches the microcanonical set $\Omega_{\varepsilon}$ (Leonarduzzi et al., 2019; Morel et al., 2023).
While having been successfully applied in a variety of domains (Bruna & Mallat, 2019; Leonarduzzi et al., 2019; Zhang & Mallat, 2021; Brochard et al., 2022; Cheng et al., 2023; Auclair et al., 2023; Morel et al., 2023), the MGDM suffers from a problem of entropy collapse. Indeed, each gradient step in (3) has the effect of reducing the entropy of the initial distribution, resulting in a final distribution which may yield a good log-likelihood (i.e., that will produce samples with characteristics similar to the training data), but which do not contain enough variability to constitute an accurate generative model. To remedy this, we propose a variant of the MGDM, called the mean-field microcanonical gradient descent model (MF-MGDM), which generates a batch of samples such that their mean energy vector satisfies the necessary constraints. In this model, the initial distribution is not so much contracted as transported through the energy space to the required point while maintaining much of its initial entropy. Experiments for known distributions show a significant reduction in KL divergence and results on experimental data indicate a better qualitative fit compared to samples from the conventional MGDM.
@@
The structure of this article is as follows. Section 2 surveys the literature on energy-based models and the MGDM in particular while Section 3 illustrates the entropy collapse observed in the MGDM. A proposed solution to this is introduced in Section 4 in the form of the MF-MGDM and numerical results supporting this algorithm are presented in Section 5. Python code to reproduce the results in this paper may be found at https://github.com/MarcusHaggbom/mf-mgdm.
@@ 2 Related Work
In energy-based methods, the goal is to find an energy function $E:\mathcal{X}\to\mathbb{R}$ such that the Gibbs density $q(x)\propto\exp(-E(x))$ approximates the true distribution well. This is the macrocanonical maximum entropy distribution (Cover & Thomas, 2006). Sampling from this distribution is typically done with MCMC methods, and is challenging but has been implemented in high dimensions, e.g. with Langevin dynamics (Du & Mordatch, 2019):
$x_{t+1}=x_{t}-\frac{1}{2}\gamma\nabla E(x_{t})+\sqrt{\gamma}\varepsilon_{t},$
which bears a resemblance with a noisy version of the MGDM when $E$ is the gradient descent loss $L$ in (2).
The microcanonical gradient descent model was introduced in (Bruna & Mallat, 2019). It has been used in a variety of applications, such as cosmology (Cheng et al., 2023; Auclair et al., 2023) and texture synthesis (Brochard et al., 2022; Zhang & Mallat, 2021). In these contexts, the model is often paired with various extensions of the scattering transform (Mallat, 2012) used as features in the energy function. The scattering transform is a composition of wavelet transforms and non-linearities, and can be seen as a convolutional neural net with predefined weights (Mallat, 2016). Apart from its use as an energy function in generative models, it has also found applications in image classification (Bruna & Mallat, 2013; Oyallon et al., 2019), audio similarity measurement (Andén et al., 2019; Lostanlen et al., 2021), molecular energy regression (Eickenberg et al., 2017), and heart beat classification (Warrick et al., 2020) among others.
In the context of finance, the MGDM has been used to generate sample paths of time series, with focus on the design of the energy function such that the approximations capture some of the typical structures that financial time series typically exhibit. In (Leonarduzzi et al., 2019), it is shown that the time-average of the second-order scattering transform encodes heavy tails, and that including also phase harmonic correlations (Mallat et al., 2019) encapsulates temporal asymmetries. An extension of this representation is known as the scattering spectrum (Morel et al., 2023), which provides a more sparse feature representation and is shown to capture multiscale properties of rough paths such as fractional Brownian motion.
Another popular feature representation for rough paths is the truncated signature (Lyons, 2014). The full signature of a path is a lossless representation up to time parametrization, and the truncation error decreases as the inverse of the factorial. In financial time series generation, this encoding has proved efficient for other generative models that involve training, such as variational autoencoders (Buehler et al., 2020) and Wasserstein GANs (Ni et al., 2021).
@@ 3 Overfitting to Target Energy
As stated, the MGDM works by pushing samples from the initial distribution towards some target energy vector which is our best approximation of the mean energy under the true distribution. It is evident that if we for some reason suspect we have a poor estimate of the mean energy, then we do not want to push our samples too close to the target, as we will then risk overfitting to the bias. What is also the case, however, is that the same principle holds even if we know the exact mean energy, as we will always lose entropy through the gradient descent process. It is therefore important to not overfit to the target energy, no matter how good the estimate is.
@@ 3.1 An Illustrative Example
As an example, take the AR(1) model with parameter $\varphi$ and conditional variance $\sigma^{2}$
@@
where $(\varepsilon_{i})_{i}$ is white noise, e.g. i.i.d. standard Gaussians. If $|\varphi|<1$, the process is stationary and has the marginal (unconditional) distribution $x_{i}\sim\mathcal{N}(0,\sigma^{2}/(1-\varphi^{2}))$. Assuming $x_{1}$ is drawn from this marginal, the likelihood is then
$p(x)$ $\propto\exp{\left\{-\frac{1}{2\sigma^{2}}\sum_{i=2}^{d}(x_{i}-\varphi x_{i-1})^{2}-\frac{1-\varphi^{2}}{2\sigma^{2}}x_{1}^{2}\right\}}$
$\approx\exp{\left\{\frac{\varphi}{\sigma^{2}}\sum_{i=2}^{d}x_{i}x_{i-1}-\frac{1+\varphi^{2}}{2\sigma^{2}}\sum_{i=1}^{d}x_{i}^{2}\right\}}.$
@@
$\Phi(x)=\left(\sum_{i=2}^{d}x_{i}x_{i-1},\;\sum_{i=1}^{d}x_{i}^{2}\right),$ (5)
and is by the Boltzmann equivalence principle asymptotically equivalent with the microcanonical approximation (Lanford, 1975) with energy function (5).
Let us now approximate the microcanonical model using MGDM. We thus have an initial measure $q_{0}$ (typically i.i.d. Gaussian with a zero mean and variance set to that of the target process) that is mapped through $T$ steps of the gradient descent procedure to some final measure $q_{T}$. Figure 1(a) illustrates how the initial distribution in the energy space $\Phi_{\#}q_{0}$ is mapped to its final distribution $\Phi_{\#}q_{T}$ after $T=100$ steps, bringing it closer to the target energy. The samples generated from the approximation will thus all have energies close to the target, but as can be seen in the pushforward of the true measure, $\Phi_{\#}p$, samples from the true distribution have a much greater variability in these statistics. If we instead stop the gradient descent after $T=36$, we obtain the distribution in Figure 1(b), where we have preserved more of the entropy, but at the cost of a lower likelihood.
This trade-off is described more in detail in the left of Figure 1(c). Here we see that although the reverse KL divergence decreases for the first 36 steps of the gradient descent, but each gradient step after this point increases the divergence. Splitting up the KL divergence into a negative entropy and a log-likelihood term in the right part of Figure 1(c), we see that although the log-likelihood continues to increase throughout the gradient descent process and eventually stabilizes, the entropy decrease at a near constant rate. It is this entropy collapse that causes the KL divergence to increase without bound, making clear the need for regularization. Note that the figure displays the outcome in the idealized setting, where $\alpha$ is equal to the true target energy $\mathbb{E}_{p}[\Phi(X)]$. In practice, this does not hold, and we also obtain unwanted overfitting to the bias $\alpha-\mathbb{E}_{p}[\Phi(X)]$.
(a) After 100 steps
(b) After 36 steps
(c) Reverse KL divergence
Figure 1: AR(1) process with $\varphi=0.1$ and $\sigma^{2}=0.99$. Top row illustrates densities of $\Phi(X)$, using fitted 2D Gaussians, at different stages of the descent.
@@ 3.2 KL Divergence
Using the reverse KL divergence allows us to analyze the method in examples like the AR(1) model where we have access to the likelihood. If the step size $\gamma$ is small enough, the gradient step (3) is typically contractive and the descent algorithm can be seen as a contractive residual flow. We can therefore write the log-likelihood $\log q_{T}$ as
$\log q_{T}(x)=\log q_{0}(z)-\sum_{t=0}^{T-1}\log|\det J_{g}(G_{t}(z))|,$
where $G_{t}$ denotes $t$ compositions of $g$ (with $G_{0}$ being identity), and $G_{T}(z)=x$. When $\Phi$ is efficiently computable, e.g. does not contain a multitude of matrix–vector multiplications with big matrices, and $d$ is not too big, this Jacobian determinant can be computed in reasonable time with automatic differentiation, e.g. using torch.func in PyTorch (Paszke et al., 2019) which allows for vectorized computation.
Going back to the AR(1) example, Figure 1(c) illustrates how the reverse KL divergence starts increasing after a certain number of gradient steps. The marginal increase in expected log-likelihood at each step eventually diminishes as the energy vector approaches the target, whereas the decrease in entropy stays consistent. The figure suggest early stopping as regularization. However, as seen in the figure, even early stopping does not produce a satisfactory distribution. Although it has a higher entropy, the energy distribution is still not centered correctly compared to the true energy distribution. As a result, while there is a trade-off between entropy and log-likelihood, it is a false trade-off in that minimizing the KL divergence leaves us with a poor entropy and a poor log-likelihood.

### v2

@@ 1 Introduction
The defining characteristic of a well-behaved generative model is the balance between its ability to, on the one hand, produce samples that are typical of the training data, while on the other hand having a significant amount of diversity within its samples. For example, a generative adversarial network (GAN) which has suffered mode collapse could produce great samples within one mode but not others. Similarly, the empirical distribution of the training data approximates the training data well but is useless for generating new samples, while a Gaussian white noise model may produce highly diverse samples that have no relation to the training data. Formally, we can view this in terms of the reverse Kullback–Leibler (KL) divergence [1] of the generative model $q$ with respect to the true distribution $p$ on the sample space $\mathcal{X}$:
$\mathcal{D}_{\text{KL}}(q\parallel p)=-H(q)-\mathbb{E}_{q}[\log p(X)],$ (1)
@@
One popular family of generative models is that of the energy-based model (EBM) [2], also known as a canonical or macrocanonical ensemble [3], typically formulated as the Gibbs or Boltzmann distribution $q(x)\propto\exp(-\beta\cdot\Phi(x))$ for a energy function $\Phi:\mathcal{X}\to\mathbb{R}^{K}$ and parameter vector $\beta\in\mathbb{R}^{K}$ (the inverse temperature). This is the distribution that maximizes the entropy $H(q)$ subject to the moment constraint $\mathbb{E}_{q}[\Phi(X)]=\alpha$ for some target energy vector $\alpha\in\mathbb{R}^{K}$ [4].
In this work, we tackle the one-shot learning problem, where we are given $\Phi$ and $\alpha=\Phi(y)$ is obtained from some observation $y\in\mathcal{X}$. Here, $\Phi$ may be given by some domain-specific design or earlier learning procedures. Using the macrocanonical approach here suffers from two main challenges, namely determining $\beta$ and sampling, both nontrivial in the general case and in particular when $\mathcal{X}$ is high-dimensional. As a remedy to the first issue is the microcanonical ensemble [5, 6, 7], which is also a maximum-entropy distribution but constrained to distributions with support in the microcanonical set of width $\varepsilon>0$,
$\Omega_{\varepsilon}\coloneqq\{x\in\mathcal{X}:\|\Phi(x)-\alpha\|\leq\varepsilon\}.$ (2)
Maximizing the entropy here implies that the distribution is uniform over this set. Thus, the entropy is equal to the log of the volume of $\Omega_{\varepsilon}$ which is increasing in $\varepsilon$. This approximation relies on the assumption that $\Phi(X)$ concentrates around its mean with high probability under the true distribution $p$, which is the case for most stationary time series of sufficiently long duration and when $\Phi$ is defined as the time average of time-shift equivariant potentials. The parameter $\varepsilon$ can then be adjusted to match this concentration of $\Phi(X)$.
@@
While the microcanonical ensemble avoids the issue of estimating $\beta$ in the macrocanonical model, sampling in high-dimensional spaces remains challenging. To mitigate this, the microcanonical gradient descent model (MGDM) was introduced by Bruna and Mallat [8] as an approximation of the microcanonical ensemble which is easier to sample from, and has been successfully applied in a variety of domains [8, 9, 10, 11, 12, 13, 14]. The MGDM is defined as the pushforward of Gaussian white noise by way of a sequence of gradient descent steps that seek to minimize the objective
$L(x)\coloneqq\frac{1}{2}\|\Phi(x)-\alpha\|^{2}.$ (3)
@@
Thus, taking $\mathcal{X}=\mathbb{R}^{d}$, samples from the MGDM are generated by sampling $x_{0}$ from $\mathcal{N}(0,\sigma^{2}\text{I}_{d})$ for some initial variance $\sigma^{2}$ and updating the sample using
$x_{t+1}=g(x_{t})\coloneqq x_{t}-\gamma J_{\Phi}^{\top}(x_{t})(\Phi(x_{t})-\alpha),$ (4)
@@
where $\gamma$ is the step size and $J_{\Phi}\in\mathbb{R}^{K\times d}$ is the Jacobian of $\Phi$. This is typically iterated for a fixed number of steps $T$ or until $x_{t}$ reaches the microcanonical set $\Omega_{\varepsilon}$ for some fixed $\varepsilon$ [9, 10].
Despite its success, MGDM can be shown to suffer from entropy collapse in many cases, resulting in a model that is able to produce typical samples but lacks sufficient variability. We shall see that this is due to the contraction of the distribution that typically occurs with each gradient step, reducing the entropy and leading to a higher KL divergence. To remedy this, we propose a new variant of the MGDM, called the mean-field microcanonical gradient descent model (MF–MGDM), which generates a batch of samples $\bm{x}\coloneqq\{x^{(n)}\}_{n=1}^{N}$ such that their mean energy vector satisfies the necessary constraints, effectively replacing $\Phi$ in (3) with the batch mean
$\overline{\Phi}(\bm{x})\coloneqq\frac{1}{N}\sum\nolimits_{n=1}^{N}\Phi(x^{(n)}).$ (5)
In this model, the initial distribution is not so much contracted as transported through the energy space to the target while maintaining more of its initial entropy. We provide a theoretical justification for this in the form of a tighter lower bound on the entropy. The resulting model combines the expressiveness of the micro- and macrocanonical ensembles with the efficient sampling of the MGDM. The choice of energy function $\Phi$ is highly dependent on the particular distribution to be approximated. To illustrate the power of the proposed approach, we therefore evaluate MF–MGDM for a range of possible functions. In each case, we see a significant improvement over the basic MGDM approach, validating the theoretical results obtained on the entropy lower bound.
@@
The structure of this article is as follows. Section 2 surveys the literature on energy-based models and the MGDM in particular, while Section 3 illustrates the entropy collapse observed in the MGDM. A proposed solution to this is introduced in Section 4 in the form of the MF–MGDM along with a lower bound on its entropy, and numerical results supporting this algorithm are presented in Section 5. Python code to reproduce the results in this paper may be found at https://github.com/MarcusHaggbom/mf-mgdm.
@@ 2 Related work
The micro- and macrocanonical ensembles are both maximum entropy distributions conditioned on the target energy $\alpha$. These are related via the Boltzmann equivalence principle [5], which states that under certain conditions of $\Phi$, they converge to the same measure as $\dim\mathcal{X}\to\infty$ and $\varepsilon\to 0$. While it is not guaranteed that a maximum entropy distribution exists in the macrocanonical case [4], the microcanonical ensemble is more general in that it allows for a wider range of energy functions [8]. Both ensembles allow sampling by MCMC methods, which is computationally challenging, but have been employed in high dimensions for EBMs [15] and score-based diffusion models [16]. This relies on sufficient mixing of the Markov chain, which is crucial for obtaining reliable Monte Carlo estimates in finite time, e.g. of the expectations in (1) when comparing models with respect to the reverse KL divergence.
The MGDM was introduced in Bruna and Mallat [8] for the purpose of facilitating sampling. Each step is deterministic, allowing us to calculate the exact likelihood of each sample, which, unlike MCMC methods, makes computing entropy comparatively easy. The MGDM has been used in a variety of applications, such as cosmology [13, 14] and texture synthesis [12, 11]. In these contexts, the model is often paired with various extensions of the scattering transform [17] used as features in the energy function. The scattering transform is a composition of wavelet transforms and non-linearities, and can be seen as a convolutional neural net with predefined weights [18]. Apart from its use as an energy function in generative models, it has also found applications in image classification [19, 20], audio similarity measurement [21, 22], molecular energy regression [23], and heart beat classification [24] among others.
In the context of finance, MGDMs coupled with variants of the scattering transform have been used to generate sample paths of time series. In Leonarduzzi et al. [9], it is shown that the time-average of the second-order scattering transform encodes heavy tails, and that including also phase harmonic correlations [25] encapsulates temporal asymmetries, both of which are typical features of financial time series. An extension of this representation is the scattering spectrum [10], which increases sparsity and better captures multiscale properties of rough paths such as fractional Brownian motion.
Another popular feature representation for rough paths is the truncated signature [26]. The full signature of a path is a lossless representation up to time parametrization, and the truncation error decreases as the inverse of the factorial of the number of included terms. Whereas the features based on the scattering transform are typically used as is, the truncated signature usually functions as a compact initial feature on top of which learning methods are applied. In financial time series generation, this encoding has proved efficient for other generative models, e.g. variational autoencoders [27] and Wasserstein GANs [28]. In principle, these learned features could serve as energy function in the canonical ensembles.
@@ 3 Overfitting to target energy
With each gradient step, the MGDM pushes the energy vector $\Phi(x)$ of a sample $x$ from the initial distribution towards the target energy $\alpha$. In doing so, however, the distribution of $x$ and $\Phi(x)$ also contracts. As a result, by the time the process reaches the microcanonical set $\Omega_{\varepsilon}$, a significant reduction of entropy has been incurred, producing a poor fit to the microcanonical ensemble.
@@ 3.1 An illustrative example
As an example, we consider the AR(1) model with parameter $\varphi$ and conditional variance $\sigma^{2}$:
@@
where $(\varepsilon_{i})_{i}$ is Gaussian white noise. If $|\varphi|<1$, the process is stationary and has the marginal distribution $x_{i}\sim\mathcal{N}(0,\sigma^{2}/(1-\varphi^{2}))$. Assuming $x_{1}$ is drawn from this marginal, the likelihood is
$p(x)$ $\propto\exp{\left\{-\frac{1}{2\sigma^{2}}\sum_{i=2}^{d}(x_{i}-\varphi x_{i-1})^{2}-\frac{1-\varphi^{2}}{2\sigma^{2}}x_{1}^{2}\right\}}\approx\exp{\left\{\frac{\varphi}{\sigma^{2}}\sum_{i=2}^{d}x_{i}x_{i-1}-\frac{1+\varphi^{2}}{2\sigma^{2}}\sum_{i=1}^{d}x_{i}^{2}\right\}}.$
@@
$\Phi(x)=\left(\frac{1}{d}\sum_{i=2}^{d}x_{i}x_{i-1},\;\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}\right),$ (7)
and is by the Boltzmann equivalence principle asymptotically equivalent with the microcanonical approximation with energy function (7).
Let us now approximate the microcanonical model using MGDM. We thus have an initial measure $q_{0}$ that is mapped through $T$ steps of gradient descent to some final measure $q_{T}$. Figure 1(a) illustrates how the initial distribution in the energy space $\Phi_{\#}q_{0}$ is mapped to its final distribution $\Phi_{\#}q_{T}$ after $T=100$ steps, bringing it close to the target energy. As can be seen in the pushforward of the true measure $\Phi_{\#}p$, however, true samples have a much greater variability in these statistics, making clear the need for regularization. If we instead stop the gradient descent earlier, after $T=36$ steps, we obtain the distribution in Figure 1(b), where we have preserved more of the entropy, but at the cost of a worse likelihood fit. The MF–MGDM, which we introduce below, performs well with respect to both aspects (Fig. 1(c)).
(a) MGDM, $T=100$
(b) MGDM, $T=36$
(c) MF–MGDM, $T=157$
Figure 1: Densities of $\Phi(X)$, using fitted 2D Gaussians, at different stages of the descent for MGDM and MF–MGDM. In (b) and (c), $T$ is the respective optimal number of steps to minimize KL divergence. The true distribution $p$ is an AR(1) process with $\varphi=0.1$ and $\sigma^{2}=0.99$.
(a) MGDM
(b) MF–MGDM
Figure 2: Reverse KL divergence for the AR(1) example. The negative entropy and expected log-likelihood are plotted on the left-hand side, and the divergence on the right.
@@ 3.2 KL divergence
Using the reverse KL divergence allows us to quantitatively analyze the method in examples like the AR(1) model where we have access to the density function of the target distribution. If $\nabla L$ is Lipschitz and the step size $\gamma$ is smaller than the Lipschitz constant, the gradient step (4) is contractive and MGDM can be seen as a contractive residual flow. The log-likelihood $\log q_{T}$ is therefore
$\begin{split}&\log q_{T}(x)=\log q_{0}(z)\\ &\quad-\sum_{t=0}^{T-1}\log|\det J_{g}(G_{t}(z))|,\end{split}$ (8)
where $G_{t}$ denotes $t$ compositions of $g$ (with $G_{0}\coloneqq\text{I}$), and $z\coloneqq G_{T}^{-1}(x)$. The Jacobian $\det J_{g}(G_{t}(z))$ is computed by automatic differentiation through torch.func in PyTorch [29, v2.1] (BSD-3). To arrive at the KL divergence, the expected values of (8) and $\log p$ in (1) are estimated by Monte Carlo.
Going back to the AR(1) example, Figure 2(a) illustrates how the reverse KL divergence attains its minimum after $T=36$ steps and then starts increasing; the improvement in likelihood fit gradually diminishes while the entropy keeps decreasing, causing an entropy collapse. In this case, the trade-off between entropy and log-likelihood is a false trade-off in that minimizing the KL divergence leaves us with a poor entropy and a poor expected log-likelihood, arguing against early stopping as a means of regularization. In contrast, we see that the proposed method, MF–MGDM, does not exhibit this problem in Figure 2(b).